OpenStudy (igreen):

Find \(a_1\) and d for for an arithmetic sequence with these terms: \(a_3 = -8\) and \(a_7 = 32\)

OpenStudy (igreen):

@hartnn @ganeshie8

OpenStudy (igreen):

@phi

OpenStudy (igreen):

..

OpenStudy (igreen):

How did you get that?

OpenStudy (anonymous):

1sec

OpenStudy (igreen):

Find the three arithmetic means in this sequence. 2, __, __, __, -18

OpenStudy (igreen):

Do you know this? @kirbykirby

OpenStudy (kirbykirby):

\(a_n=a_m+(n-m)d\) where \(a_7=a_n\) and \(a_3=a_m\)

OpenStudy (igreen):

So what am I supposed to do with that..?

OpenStudy (kirbykirby):

you can compare the indices, and see that \(n = 7\) and \(m=3\) now you can plug in those values into the formula \(32 = -8(7-3)d\)

OpenStudy (igreen):

Oh..

OpenStudy (igreen):

\(32 = -8(7 - 3)d\) \(32 = -56d + 24d\) Is that right..?

OpenStudy (kirbykirby):

yes

OpenStudy (igreen):

\(32 = -32d\) \(d = -1\)

OpenStudy (kirbykirby):

oh wait I forgot the + when I re-wrote it with numbers.. sorry: \(32 = -8\color{red}{+}(7-3)d\)

OpenStudy (igreen):

\(32 = -8 + 7d - 3d\) \(32 = -8 + 4d\) \(40 = 4\) \(d = 10\)

OpenStudy (igreen):

*4d

OpenStudy (kirbykirby):

yes that is good :) sorry again..

OpenStudy (igreen):

Lol, fine. I have a few more..I'll tag you. :)

OpenStudy (kirbykirby):

so finding \(a_1\) will be the same as above: \(\large a_{\color{red}{3}} = a_\color{blue}{1}+(\color{red}{3}-\color{blue}{1})(10)\)